OpenStudy (anonymous):

Using the Law of Cosines how would I find an angle of the triangle if a= 32, b= 37, c= 24? Explain please! :)

OpenStudy (anonymous):

I will give a medal, and a fan, but I have to go, so it may take like an hour, and I'll eventually get what you guys post :P

Parth (parthkohli):

All the angles?\[c^2 = a^2 + b^2 - 2ab\cos(\gamma)\]\[b^2 = a^2 + c^2 - 2ac \cos(\beta)\]\[a^2 = b^2 + c^2 - 2bc\cos(\alpha)\]Use the values of a,b, and c

OpenStudy (phi):

I would draw the triangle, label all the sides. then label the angle you want to solve then use the formula.

OpenStudy (anonymous):

Thank you! :)